OpenStudy (anonymous):

Y=2x^2-32 Find the x-intercepts of this equation how do u do that ?

OpenStudy (anonymous):

Help some one :(

OpenStudy (sburchette):

x-intercepts of a function occur at values of x that cause y to equal zero. To find x-intercepts, substitute 0 for y and solve for x. \[0=2x ^{2}-32 \] \[32=2x ^{2}\] (Add 32 to both sides) \[x^2=16 \] (divide both sides by 2) \[\sqrt{x^2}=\pm \sqrt{16}\] (Take the square root of both sides) \[x=\pm4\] Therefore, the x-intercepts of the function are 4 and -4. If you graph the function, you should see the curve intersect the x-axis at those values. If you substitute 4 or -4 in for x, you should get 0.